### v1

@@ 4.2 Educational Programs
5 Theory of the Double Bracketing
So far, we have observed that the double bracketing relationship, ${\theta_{\text{ATT}}^{\text{M}}}\leq{\theta_{\text{ATT}}^{\text{DIDM}}}\leq{\theta_{\text{ATT}}^{\text{DID}}}$, appears to be more of a rule than a mere coincidence in empirical studies on educational and job training programs. In this section, and the subsequent ones, we are going to provide formal theoretical explanations for this intriguing phenomenon. To ensure generality, we adopt a non-parametric and model-free framework.
@@ 5 Theory of the Double Bracketing
Each of the three components, (i)–(iii), of Assumption 5.1 is empirically testable, as it does not involve unobserved latent variables such as $Y_{t}(d)$. Furthermore, they are plausible and align with the assumptions made in the literature (e.g., Angrist and Pischke,, 2009; Ding and Li,, 2019). Specifically, part (i) requires the so-called “negative selection” that individuals who opt out from treatment (i.e., those with $W=0$) tend to have weakly higher pre-treatment (potential) outcomes $Y_{0}=Y_{0}(0)$ without treatment on average given $Y_{-s}=y$. Similarly, part (ii) requires the negative selection that individuals who opt out from treatment (i.e., those with $W=0$) tend to have no lower pre-treatment (potential) outcome $Y_{-s}=Y_{-s}(0)$ without treatment. Part (iii) requires the time series of $\{Y_{t}\}_{t}$ of outcomes for those with $W=0$ (i.e., $\{Y_{t}\}_{t}=\{Y_{t}(0)\}_{t}$) is non-explosive. In the special case where $s=0$, this requirement essentially reflects the condition that the root of the AR(1) process is less than one.
@@
6 Intuition of the Double Bracketing
Section 5 presented a formal theoretical explanation of the double bracketing relationship ${\theta_{\text{ATT}}^{\text{M}}}\leq{\theta_{\text{ATT}}^{\text{DIDM}}}\leq{\theta_{\text{ATT}}^{\text{DID}}}$ without imposing any parametric model specification. The current section develops intuition for this formal result by focusing on a simple parametric model.
Suppose the true data-generating process takes the form:
$Y_{t}=\alpha+\beta W\mathbbm{1}\{t\geq 1\}+\gamma W+\delta_{t}+\rho Y_{t-1}+\epsilon_{t},$ (1)
where the error term satisfies $E[\epsilon_{1}|Y_{-1},W]=E[\epsilon_{0}|Y_{-1},W]=0$. In this context, $\gamma$ and $\delta_{t}$ serve as two-way fixed effects and $\rho$ serves as a persistence parameter.
Note that the individual fixed effect $\gamma$ can be expressed as $\gamma=E[Y_{0}|W=1,Y_{-1}]-E[Y_{0}|W=0,Y_{-1}]$ under (1). This value, $\gamma$, is non-positive if individuals with weakly lower $Y_{0}$ tend to select into treatment $W=1$ conditional on $Y_{-1}$ as is plausibly the case with job training programs. We postulate this negative selection assumption:
[…]
We also postulate a similar negative selection assumption for $Y_{-1}$:
[…]
Note that these three conditions (2)–(4) are analogous to the three parts, (i)–(iii), of Assumption 5.1, and also in line with the assumptions invoked by Angrist and Pischke, (2009, Section 5) in developing their bracketing relationship, LDV $\leq$ FE.
${\theta_{\text{ATT}}^{\text{M}}}$ $=\beta+(1+\rho)\gamma,$ (5)
${\theta_{\text{ATT}}^{\text{DID}}}$ $=\beta+\rho\gamma+\rho(1-\rho)(E[Y_{-1}|W=0]-E[Y_{-1}|W=1]).$ (7)
${\theta_{\text{ATT}}^{\text{DID}}}-{\theta_{\text{ATT}}^{\text{DIDM}}}=\rho(1-\rho)(E[Y_{-1}|W=0]-E[Y_{-1}|W=1])\geq 0,$ (9)
where the inequality follows from $E[Y_{-1}|W=0]-E[Y_{-1}|W=1]\geq 0$ due to (3) and $\rho\in[0,1]$ due to (4).
Combining (8)–(9) yields the double bracketing relationship,
which holds true under the simple parametric model (1) with the plausible restrictions (2)–(4) motivated by Angrist and Pischke, (2009, Section 5). In this manner, the simple parametric model (1) allows for an intuitive and tractable illustration of the general double bracketing result presented in Section 5.
@@ Appendix C Detailed Calculations for Section 6
This appendix section presents detailed calculations to derive the expressions (5)–(7) in Section 6.

### v2

@@ 3.1 The M, DID, and DIDM Estimands
3.2 Mutual Non-Nestedness of The M, DID, and DIDM Conditions
The three identifying assumptions underlying the estimands ${\theta_{\text{ATT}}^{\text{M}}}$, ${\theta_{\text{ATT}}^{\text{DID}}}$, and ${\theta_{\text{ATT}}^{\text{DIDM}}}$ each restrict the data‐generating process (DGP) in a distinct way. No one assumption nests another; that is, none of the conditions implies or is implied by any of the others. In the following, we provide intuitive counterexamples—some leveraging differences in sample composition—to illustrate these distinctions.
3.2.1 Condition M Holds but Condition DIDM Fails
Consider the following data-generating process (DGP) for the potential outcomes:
$Y_{0}(0)$ $=\mu(W)+\eta_{0}(0),$ $Y_{1}(0)$ $=\eta_{1}(0),$
$Y_{0}(1)$ $=\mu(W)+\eta_{0}(1),$ $Y_{1}(1)$ $=\eta_{1}(1),$
where $(\eta_{0}(0),\eta_{0}(1),\eta_{1}(0),\eta_{1}(1))\perp\!\!\!\perp(W,Y_{-s})$ and $\mu(\cdot)$ is an arbitrary function such that $\mu(0)\neq\mu(1)$. Under this DGP, $(Y_{1}(1),Y_{1}(0))\perp\!\!\!\perp W|Y_{-s}$ holds, but ${\text{E}}[Y_{1}(0)-Y_{0}(0)|Y_{-s},W=0]={\text{E}}[\eta_{1}(0)-\eta_{0}(0)]-\mu(0)\neq{\text{E}}[\eta_{1}(0)-\eta_{0}(0)]-\mu(1)={\text{E}}[Y_{1}(0)-Y_{0}(0)|Y_{-s},W=1]$. In other words, Condition M holds, but Condition DIDM fails.
3.2.2 Condition DIDM Holds but Condition M Fails
Conversely, consider the following DGP for the potential outcomes:
$Y_{0}(0)$ $=\mu(W)+\eta_{0}(0),$ $Y_{1}(0)$ $=\mu(W)+\eta_{1}(0),$
$Y_{0}(1)$ $=\mu(W)+\eta_{0}(1),$ $Y_{1}(1)$ $=\mu(W)+\eta_{1}(1),$
where $(\eta_{0}(0),\eta_{0}(1),\eta_{1}(0),\eta_{1}(1))\perp\!\!\!\perp(W,Y_{-s})$ and $\mu(\cdot)$ is an arbitrary function such that $\mu(0)\neq\mu(1)$. Under this DGP, ${\text{E}}[Y_{1}(w)|Y_{-s},W=0]=\mu(0)+{\text{E}}[\eta_{1}(0)]\neq\mu(1)+{\text{E}}[\eta_{1}(0)]={\text{E}}[Y_{1}(w)|Y_{-s},W=1]$, and hence, $(Y_{1}(1),Y_{1}(0))\not\perp\!\!\!\perp W|Y_{-s}$, but ${\text{E}}[Y_{1}(0)-Y_{0}(0)|Y_{-s},W=0]={\text{E}}[\eta_{1}(0)-\eta_{0}(0)]={\text{E}}[Y_{1}(0)-Y_{0}(0)|Y_{-s},W=1]$ holds. In other words, Condition M fails, but Condition DIDM holds.
3.2.3 Condition DIDM Holds but Condition DID Fails
Consider the following DGP for untreated potential outcomes:
$Y_{0}(0)$ $=\mu(Y_{-1})+\eta_{0}(0),$ $Y_{1}(0)$ $=\mu(Y_{-1})+\eta_{1}(0),$
where $E[\eta_{1}(0)-\eta_{0}(0)|Y_{-1},W]=Y_{-1}$ and $E[Y_{-1}|W=0]\neq E[Y_{-1}|W=1]$. Under this DGP, we have $E[Y_{1}(0)-Y_{0}(0)|Y_{-1},W=0]=Y_{-1}=E[Y_{1}(0)-Y_{0}(0)|Y_{-1},W=1]$, but $E[Y_{1}(0)-Y_{0}(0)|W=0]=E[E[\eta_{1}(0)-\eta_{0}(0)|Y_{-1},W=0]|W=0]=E[Y_{-1}|W=0]\neq E[Y_{-1}|W=1]=E[E[\eta_{1}(0)-\eta_{0}(0)|Y_{-1},W=1]|W=1]=E[Y_{1}(0)-Y_{0}(0)|W=1].$ In other words, Condition DIDM holds, but Condition DID fails. Intuitively, even if the parallel trends hold at every level of $Y_{-1}$, differences in the distribution of $Y_{-1}$ between treatment groups can lead to unequal unconditional trends.
3.2.4 Condition DID Holds but Condition DIDM Fails
Conversely, consider the following DGP for untreated potential outcomes:
$Y_{0}(0)$ $=\mu(Y_{-1})+\eta_{0}(0),$ $Y_{1}(0)$ $=\mu(Y_{-1})+\eta_{1}(0),$
where $E[\eta_{1}(0)-\eta_{0}(0)|Y_{-1},W]=(2W-1)Y_{-1}$ and $E[Y_{-1}|W=0]+E[Y_{-1}|W=1]=0$. Under this DGP, $E[Y_{1}(0)-Y_{0}(0)|W=0]=E[E[Y_{1}(0)-Y_{0}(0)|Y_{-1},W=0]|W=0]=E[-Y_{-1}|W=0]=E[Y_{-1}|W=1]=E[E[Y_{1}(0)-Y_{0}(0)|Y_{-1},W=1]|W=1]=E[Y_{1}(0)-Y_{0}(0)|W=1],$ but $E[Y_{1}(0)-Y_{0}(0)|Y_{-1},W=0]=-Y_{-1}\neq Y_{-1}=E[Y_{1}(0)-Y_{0}(0)|Y_{-1},W=1].$ In other words, Condition DID holds but Condition DIDM fails. Intuitively, even if the overall (unconditional) trends are equal (due to cancellation when aggregating over $Y_{-1}$), the trends might differ for each subpopulation defined by $Y_{-1}$.
Implication for Practice.
These examples underscore that each assumption addresses a different facet of the DGP. Practitioners must ex ante commit to one of these assumptions based on the context and the nature of available data. Whether one relies on matching (Condition M), unconditional parallel trends (DID), or conditional parallel trends (DIDM) is not a matter of nested robustness but of fundamentally different identifying restrictions, each carrying its own trade-offs and implications for causal inference.
@@ 4.2 Educational Programs
5 The Double Bracketing: A Parametric Approach
So far, we have observed that the double bracketing relationship, ${\theta_{\text{ATT}}^{\text{M}}}\leq{\theta_{\text{ATT}}^{\text{DIDM}}}\leq{\theta_{\text{ATT}}^{\text{DID}}}$, appears to be more of a rule than a mere coincidence in empirical studies on educational and job training programs. In this section, and the two subsequent ones, we are going to provide theoretical explanations for this intriguing phenomenon.
While general theories will be presented in the subsequent sections, we begin by illustrating how the double bracketing relationship, ${\theta_{\text{ATT}}^{\text{M}}}\leq{\theta_{\text{ATT}}^{\text{DIDM}}}\leq{\theta_{\text{ATT}}^{\text{DID}}}$ arises under a familiar parametric setting. In particular, suppose the true data-generating process takes the form:
$Y_{i,t}=\alpha_{i}+\beta W_{i}\mathbbm{1}\{t\geq 1\}+\gamma W_{i}+\delta_{t}+\rho Y_{i,t-1}+\epsilon_{i,t},$ (1)
where the error term satisfies
$E[\epsilon_{i,1}|Y_{i,-1},W_{i}]=E[\epsilon_{i,0}|Y_{i,-1},W_{i}]=0.$ (2)
In this context, $\alpha_{i}$ and $\delta_{t}$ serve as two-way fixed effects and $\rho$ serves as a persistence parameter.
Note that $\gamma$ can be expressed as $\gamma=E[Y_{i,0}|W_{i}=1,Y_{i,-1}]-E[Y_{i,0}|W_{i}=0,Y_{i,-1}]$ under (1)–(2). This value, $\gamma$, is non-positive if individuals with weakly lower $Y_{i,0}$ tend to select into treatment $W_{i}=1$ conditional on $Y_{i,-1}$ as is plausibly the case with job training programs. We postulate this negative selection assumption:
[…]
We also postulate a similar negative selection assumption for $Y_{i,-1}$:
[…]
Note that these three conditions (3)–(5) are analogous to the assumptions invoked by Angrist and Pischke, (2009, Section 5) in developing their bracketing relationship, LDV $\leq$ FE.
${\theta_{\text{ATT}}^{\text{M}}}$ $=\beta+(1+\rho)\gamma,$ (6)
${\theta_{\text{ATT}}^{\text{DID}}}$ $=\beta+\rho\gamma+\rho(1-\rho)(E[Y_{i,-1}|W_{i}=0]-E[Y_{i,-1}|W_{i}=1]).$ (8)
${\theta_{\text{ATT}}^{\text{DID}}}-{\theta_{\text{ATT}}^{\text{DIDM}}}=\rho(1-\rho)(E[Y_{i,-1}|W_{i}=0]-E[Y_{i,-1}|W_{i}=1])\geq 0,$ (10)
where the inequality follows from $E[Y_{i,-1}|W_{i}=0]-E[Y_{i,-1}|W_{i}=1]\geq 0$ due to (4) and $\rho\in[0,1]$ due to (5).
Combining (9)–(10) yields the double bracketing relationship,
which holds true under the simple parametric model (1) with the plausible restrictions (3)–(5) motivated by Angrist and Pischke, (2009, Section 5).
6 The Double Bracketing in Nonparametric Frameworks
The previous section derived the double-bracketing relationship ${\theta_{\text{ATT}}^{\text{M}}}\leq{\theta_{\text{ATT}}^{\text{DIDM}}}\leq{\theta_{\text{ATT}}^{\text{DID}}}$ under a simple parametric model framework. To ensure generality, we adopt a non-parametric and model-free framework in the current section.
@@ 6 The Double Bracketing in Nonparametric Frameworks
The three components, (i)–(iii), of Assumption 6.1 are analogous to (3)–(5), respectively. Each of the three components is empirically testable, as it does not involve unobserved latent variables such as $Y_{t}(d)$. Furthermore, they are plausible and align with the assumptions made in the literature (e.g., Angrist and Pischke,, 2009; Ding and Li,, 2019). Specifically, part (i) requires the so-called “negative selection” that individuals who opt out from treatment (i.e., those with $W=0$) tend to have weakly higher pre-treatment (potential) outcomes $Y_{0}=Y_{0}(0)$ without treatment on average given $Y_{-s}=y$. Similarly, part (ii) requires the negative selection that individuals who opt out from treatment (i.e., those with $W=0$) tend to have no lower pre-treatment (potential) outcome $Y_{-s}=Y_{-s}(0)$ without treatment. Part (iii) requires the time series $\{Y_{t}\}_{t}$ of outcomes for those with $W=0$ (i.e., $\{Y_{t}\}_{t}=\{Y_{t}(0)\}_{t}$) is stable over time. In real-world settings, especially among populations with limited socioeconomic status (SES), it may be reasonable to expect that outcomes do not escalate dramatically without intervention. For example, in educational programs targeting low-SES individuals, we wouldn’t anticipate rapid, unsustainable improvements in outcomes without structured support on average. In the special case where $s=0$, this requirement essentially reflects the condition that the root of the AR(1) process is less than one, which is equivalent to stationarity, as deployed in, e.g., Chetty et al., 2014a .
@@ Appendix C Detailed Calculations for Section 5
This appendix section presents detailed calculations to derive the expressions (6)–(8) in Section 5. We omit the $i$ subscript throughout this appendix section.